OpenStudy (anonymous):

Solve for x-x^3=2x^4+3x^2-5

OpenStudy (anonymous):

Anybody? I seriously cant do this..

OpenStudy (jack1):

1. add like terms... there are none 2. put in terms of LHS = 0 3. factorize 4. solve

OpenStudy (anonymous):

Here are the options: A)X=-.87, X=1.16 B)X=-1.56, X=1 C)X=+-1 D)X=0

OpenStudy (jack1):

step 1 . done step 2. -2x^4 - x^3 - 3x^2 + x + 5 = 0 follow so far dude?

OpenStudy (anonymous):

Yep, i think so.

OpenStudy (jack1):

step 3. - ((x - 1) (x + 1) (2x^2 + x + 5)) = 0 which is the same as writing ((x - 1) (x + 1) (2x^2 + x + 5)) = 0 (ie if u multiply both sides of the equals sign by -1 = above answer so as (x - 1) (x + 1) (2x^2 + x + 5) = 0 either (x - 1) = 0 or (x + 1) = 0 or (2x^2 + x + 5) = 0

OpenStudy (jack1):

now at your level you only need to be focusing on real solutions, not complex/ imaginary numbers, so you can ignore the line: or (2x^2 + x + 5) = 0 so your answers are either: (x - 1) = 0, so x must = +1 (x + 1) = 0, so x must equal -1

OpenStudy (jack1):

all good dude?

OpenStudy (anonymous):

yes, thanks so much!

OpenStudy (jack1):

always welcome man, slaters